# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

### v1

@@ 3.1 Hindsight Mitigates Misalignment
The foresight value at time $k\geq 0$ is the expected reward-to-go from $k$ to $k+N$ under the human’s subjective prediction of future outcomes. Foresight uses the human’s own internal predictive model of what states and actions might arise. Formally:
$\overrightarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{\begin{subarray}{c}{s}_{t}\sim P(\cdot\mid{z}^{\textit{H}}_{t}),~{}{a}^{\textit{H}}_{t}\sim{\pi}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{t})\end{subarray}}\left[\sum_{t=k}^{k+N}\gamma^{t}r({s}_{t},{a}^{\textit{H}}_{t})\right],$ (3)
where $P(\cdot\mid{z}^{\textit{H}}_{t})$ is the human’s internal predictive distribution over future states. Neither the states ${s}_{t}$ nor the rewards here are necessarily generated by the environment’s true transition kernel; rather, they are drawn from the human’s imaginative predictions.
Figure 2: Illustration of hindsight’s advantage: Delaying human feedback until the human has experienced the outcome corresponding to the bulk of reward significantly mitigates the misalignment in the AI’s learned reward model.
Using this foresight-based feedback, the human evaluator relies on predictions of future actions and states, which can be influenced by the AI. Consequently, the predicted human utility using foresight is: $\tilde{{U}}^{\textit{H}}_{T+N}({\pi}^{\textit{AI}})\;=\;\overrightarrow{{V}}_{T+N}({z}^{\textit{H}}_{T}).$ However, RL from Hindsight Simulation (RLHS) aims to reduce the human’s reliance on such predictions by shifting the evaluation from foresight to hindsight value. By simulating the downstream acting phase for $N$ steps, the human can directly observe realized outcomes rather than guess them, thus: ${U}^{\textit{H}}_{T+N}({\pi}^{\textit{AI}})\;=\;\overleftarrow{{V}}_{T+N}({z}^{\textit{H}}_{T}).$
In hindsight-based evaluation, the states $\{{s}_{t}\}$ come from actual world rollouts rather than predictions conditioned on the human’s internal state which can be influenced by the AI. This makes hindsight feedback substantially more robust and less prone to reward hacking, as it depends on concrete observed outcomes rather than speculative future scenarios influenced by the AI’s advice.
In the following, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility. Consider an oracle aligned AI policy ${\pi}^{*}$ that operates knowing the human preference ${\theta}^{\textit{H}}$. The following lemma establishes that, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$, the difference in finite-hindsight utility estimation becomes an exponentially accurate estimate of the difference in true utility as the hindsight horizon $N$ increases.
Lemma 1.
${U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})-{U}_{N}^{\textit{H}}(\tilde{\pi}^{\textit{AI}})\in\mathcal{B}\Big{(}{U}^{\textit{H}}({\pi}^{\textit{AI}})-{U}^{\textit{H}}(\tilde{\pi}^{\textit{AI}}),\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\,.$
Proof.
Theorem 1.
Suppose the human evaluator is presented a finite-horizon hindsight of $N$ steps and makes Boltzmann-rational binary preference choices with inverse temperature $\beta$. Then the probability that the human prefers a hindsight observation ${\mathbf{{o}}}_{0:T+N}$ over another $\bar{\mathbf{{o}}}_{0:T+N}$ from the same initial information state $P({\mathbf{{o}}}_{0:T+N}\succ\bar{\mathbf{{o}}}_{0:T+N})$ is within the range
3.2 Implementation: Hindsight Simulation with AI Feedback
Figure 3: Qualitative results for Llama-2-7b trained with either immediate feedback (RLHF) or partial hindsight (RLHS). The RLHF model (trained with immediate feedback) deceives the user by falsely claiming Options A and C meet the customer’s 8K resolution requirement, though neither does. In contrast, the RLHS model truthfully states that none of the options include 8K resolution.
@@ 3.2 Implementation: Hindsight Simulation with AI Feedback
Hindsight Simulation. While we have demonstrated theoretically that providing hindsight can mitigate misalignment in RLHF, exposing humans to real consequences can circumvent material and ethical difficulties. To address this, we introduce the concept of hindsight simulation—the namesake of our core contribution, RLHS —which allows evaluators, whether human or AI, to make more informed decisions based on simulated outcomes. In practice, hindsight simulation can involve collecting feedback from human evaluators or employing another language model as a proxy to simulate the feedback process. After an evaluator makes a decision based on their interaction with the AI (e.g., purchasing an item), the system provides ground truth information about the outcome, i.e., the hindsight (e.g., whether the purchased item meets the desired criteria). The evaluator then provides feedback informed by both the decision’s outcome and their prior interaction with the model.
This feedback typically takes the form of a rating or binary preference, similar to the feedback used in conventional RLHF. However, unlike the immediate feedback provided solely during an interaction without access to the decision’s consequences, feedback obtained through hindsight simulation is more informed as it incorporates long-term outcomes. This aligns with the reasoning presented in Section 3.1 and demonstrates the potential for improving alignment through simulated hindsight.
@@
We implement this approach with two subroutines: (i) partial hindsight, where only a limited set of hindsight information is available to the agent, in a way that more closely matches real-world scenarios, and (ii) oracle hindsight, where the agent has access to full set of hindsight information. In our empirical studies employing both partial and oracle hindsight, we provide insights into how extending the hindsight step (i.e., revealing additional outcome information to the agent) can improve the alignment performance of the model.
@@
Illustrative Example: Marketplace Chatbot. We demonstrate the practical impact of RLHS by applying it to a marketplace AI chatbot. The chatbot’s goal is to assist customers in making purchasing decisions by providing recommendations based on available product information. We assume that both customers and the chatbot have access to some public information, such as a list of items and their prices, but customers have their internal preferences, e.g., wanting a TV with 8K resolution, that are unknown to the chatbot. To the best of our knowledge, existing RLHF schemes deployed for training marketplace chatbots (e.g., Amazon, 2024) use customer feedback solely based on the interaction (i.e., if they feel happy about the chatbot’s service) but not on the outcome (i.e., if the purchased item actually meets their preferences), potentially causing misalignment.
Our proposed hindsight simulation approach aims to mitigate this issue by deferring the humans’ feedback until they have been informed of the outcome of their decisions, e.g., they have received the product and verified that their expectations are (not) met. In hindsight simulation, the simulated customer interacts with the chatbot, makes a purchasing decision, checks the outcome (hindsight) provided by the system, and provides feedback on the customer’s satisfaction with the service.
@@ 4.1 Data Collection
Preference Data Collection. Our training data collection process closely follows the standard RLHF data collection pipeline (Stiennon et al., 2020; Ouyang et al., 2022), where feedback is collected based on comparisons between outputs. However, instead of relying on real human feedback, we employed a strong large language model (LLM) model as a judge to simulate human interactions with the chatbot and provide feedback. For real-world online marketplace chatbots like the Amazon Rufus (Amazon, 2024), human feedback is typically given as a rating at the end of the interaction. However, human users tend to compare their current experience with previous ones when assigning ratings. To capture this behavior, we simulate users comparing services from two different stores and selecting their preferred option, rather than rating each scenario in isolation. This closely aligns with the preference-based data collection method used in prior work (Stiennon et al., 2020; Ouyang et al., 2022), where users provide feedback by comparing responses instead of giving individual ratings.
Decision-making simulation. While collecting the preference data, our simulated human (strong model) takes on three roles: interacting with the chatbot, making decisions, and providing feedback. To ensure accurate decision-making and feedback, we adapted the approach in introspective planning (Liang et al., ). First, we frame the decision-making problem as a multiple-choice question with four options: (A) Buy option A, (B) Buy option B, (C) Buy option C, or (D) Do not buy anything. We then ask the LLMs to perform Chain-of-Thought reasoning (Wei et al., 2022), querying the next token probabilities to select the best option from ${A,B,C,D}$. This approach can reduce the language agent’s uncertainty. We apply a similar method for comparing services between two stores.
Dataset Details. In our experiments, we used both Llama-2-7B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024) as the AI assistants, and Llama-3.1-70B (Dubey et al., 2024) as the simulated human to interact with the AI assistant and provide feedback. We collected 11,000 preference data points for each AI assistant model, with 10,000 used for training and 1,000 for validation. We also generated a test set of 1,200 examples to evaluate performance across different customer scenarios.
@@ 4.2 Experiment Setup
Environment Details. In each of our simulated marketplace scenarios there are 10 candidate items, each characterized by 8 features and a price. Each feature can be categorized in two ways: (1) The item either has or lacks a specific feature (e.g., a TV with HDR vs. without HDR), and (2) The feature can vary in types (e.g., 8K resolution vs. 4K resolution). While in most cases the chatbot has access to this information, there are instances where it is uncertain about a particular feature (e.g., resolution not specified). We will examine these scenarios and investigate when and how the AI acts deceptively.
@@
The satisfaction rating reflects the user’s evaluation of the chatbot’s service, measured on a 5-point Likert scale ranging from 1 (very dissatisfied) to 5 (very satisfied). For the experimental results shown in Fig. 4 and Fig. 5, these ratings were normalized to a scale between -1 and 1, which ensure that the true utility and satisfaction ratings are on the same scale for clearer comparison. Additional results using the original Likert scale are provided in Appendix A. Furthermore, we quantified two metrics in the human study: regret rate, which measures how often users regret their decisions, and hallucination rate, which measures how truthful the language model is.
Training algorithms. We explored both online and offline preference optimization methods to align our language model with human preferences. In our online approach, we trained a reward model on the preference data. The language model then interacted with the environment by generating responses and receiving reward signals from this reward model. We utilized Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tuned the model iteratively to maximize these rewards. For the offline approach, we experimented with Direct Preference Optimization (DPO) (Rafailov et al., 2024), which aligns language models with human preferences without the need for an explicit reward model. We apply LoRA fine-tuning (Hu et al., 2021) for both algorithms to efficiently update model parameters. Further details of these methods are included in Appendix B.
Figure 4: Results on Llama-2-7b trained with PPO. Left: Demonstrates the Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Shows how partial hindsight mitigate the misalignment. Right: Shows the alignment achieved with oracle hindsight.
@@ 5 Simulation Results
Misalignment between satisfaction rating and real utility. When using standard RLHF (Ouyang et al., 2022), we observe significant misalignment between user satisfaction ratings and true utility as training progresses (left plot in Figs. 4 and 5). While the satisfaction rating steadily increases, indicating that the language model is learning to deliver responses that receive higher immediate user approval, the true utility shows a sharp decline. This suggests that while the chatbot’s responses may appear more polished or helpful in the short term, they are in fact becoming less aligned with the user’s true needs or long-term goals. As a result, while users may initially perceive the chatbot’s responses as helpful, they are frequently misled and ultimately dissatisfied with their final outcomes. This highlights a fundamental flaw in using standard RLHF with immediate feedback, as it risks optimizing for superficial satisfaction at the expense of true utility.
Metric DPO PPO
IF PHS OHS IF PHS OHS
Rating $\uparrow$ $0.95_{\pm 0.028}$ $0.35_{\pm 0.032}$ $0.33_{\pm 0.036}$ $0.97_{\pm 0.021}$ $0.41_{\pm 0.026}$ $0.31_{\pm 0.024}$
True Utility $\uparrow$ $-0.51_{\pm 0.03}$ $0.18_{\pm 0.023}$ $0.23_{\pm 0.026}$ $-0.71_{\pm 0.029}$ $0.18_{\pm 0.025}$ $0.24_{\pm 0.031}$
Table 1: Comparison of performance metrics (Rating and True Utility) across models trained with DPO and PPO under three feedback conditions: Immediate Feedback (IF), Partial Hindsight Simulation (PHS), and Oracle Hindsight Simulation (OHS). Ratings are higher when trained with immediate feedback but lead to lower real utility, indicating potential misalignment between perceived satisfaction and actual utility. Hindsight simulations significantly improve the true utility.
Hindsight simulation effectively mitigates misalignment. As shown in Fig. 4 (left), relying on immediate feedback leads to a steady decline in real utility, ultimately resulting in negative overall utility. In contrast, hindsight simulation consistently improves utility throughout training, eventually achieving positive utility, as in Fig. 4 (middle). It aligns upward trends in both real utility and satisfaction ratings, significantly reducing the gap between them. The qualitative results shown in Fig. 3 further support our claim. When the AI assistant is trained on immediate feedback, it deceptively claims that both Options A and C meet the requirements of the (simulated) customer for 8K resolution, though neither actually does. In contrast, training with partial hindsight leads to truthful responses, acknowledging that none of the options meet the 8K resolution requirement. This shows that while traditional RLHF with immediate feedback may cause misalignment, hindsight simulation mitigates this issue, improving the overall helpfulness and honesty of language agents.
@@ 6 Human Study
Our human study had two goals: (Goal 1) evaluate the performance of models trained with immediate feedback vs. hindsight simulation, (Goal 2) assess how hindsight information affects user satisfaction. To achieve these goals, we designed two similar human experiments. Both experiments used Llama-3-8b (Dubey et al., 2024) trained with DPO using either immediate feedback or partial hindsight. We conducted online human experiments via Prolific (Palan & Schitter, 2018), involving 200 participants across 10 scenarios, randomly sampled from a test set of 1,200. For each scenario, 20 participants were randomly assigned to one of two conditions: 10 interacting with the RLHF model and 10 with the RLHS model. We report specific details for participant recruitment, compensation, and IRB approval in Appendix D.2.
@@
Hindsight information is then introduced. Buyers learn whether the item meets their requirements (e.g., whether the item has the desired feature) while non-buyers receive no additional information. Participants then provide a second satisfaction rating, referred to as the hindsight rating, which evaluates their long-term satisfaction after considering the hindsight information. This step allows us to assess the impact of hindsight information on user satisfaction (Goal 2). Finally, buyers may keep or return the item, enabling us to quantify the regret rate.
Statistical Hypothesis Testing. We conducted experiments to test four hypotheses, using one-tailed and standard t-tests for the first three hypotheses (Fisher, 1970), and Pearson’s correlation coefficient for the fourth (Sedgwick, 2012). The one-tailed t-test used for Hypotheses 1, 2, and 3 is outlined below. The null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{1}$) are defined as:
[…]
Here, $\mu_{1}$ and $\mu_{2}$ represent the mean satisfaction for Group 1 and Group 2, respectively. The two-tailed t-test follows a similar format but tests for any significant difference between the group means.
@@
We evaluated hindsight ratings for two models: Group 1 (RLHS) and Group 2 (RLHF). The hypothesis test resulted in $p=4\times 10^{-8}$, well below the significance threshold of 0.001. When reversing the groups for regret rates, the test yielded $p=5\times 10^{-5}$ again below 0.001.
Hypothesis 2: Models trained with RLHF using immediate feedback often experience a notable decline in user satisfaction once future outcomes are revealed, and RLHS mitigates this decline.
Group 1 consisted of users interacting with RLHF without hindsight feedback, and Group 2 received hindsight feedback. The hypothesis test gave $p=4\times 10^{-9}$, confirming a significant decline in user satisfaction. To demonstrate that RLHS mitigates this decline, we ran a two-tailed t-test comparing immediate and hindsight ratings, and find that this decline is likely no longer present ($p=0.90$).
Hypothesis 3: RLHS leads to significantly higher true utility than RLHF.
@@
Hypothesis 4: Models trained with RLHS are more truthful, presenting a strong correlation between their high immediate user satisfaction rate (subjective) and high true utility (objective).
@@
Model Immediate rating Hindsight rating True utility Regret rate
@@
Table 2: Performance comparison between RLHF and RLHS models across multiple metrics. While RLHF shows higher immediate satisfaction, RLHS outperforms in hindsight rating, true utility, and regret rate, indicating better long-term alignment with user preferences and reduced regret.

### v2

@@ 3.1 Hindsight Mitigates Misalignment
The human’s foresight value at time $k\geq 0$ is the utility expected from $k$ to $k+N$ under the human’s subjective prediction of future outcomes:
$\overrightarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{s_{k}\sim b^{\textit{H}}_{k}(\cdot|{z}^{\textit{H}}_{k})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k})\bigr{]}$ (3)
where $b_{k}^{\textit{H}}(\cdot\mid{z}^{\textit{H}}_{k})$ is the human’s belief over the state, informed by their internal state after the interaction.
The above expectation is taken over the human’s belief. Whether we consider the belief of the user or a separate evaluator, it is a function of the internal state ${z}^{\textit{H}}_{k}$ resulting from the interaction, and therefore it can, in general, be influenced by the AI output.
Definition 4 (Expected Hindsight Value).
The AI’s expected hindsight value at time $k\geq 0$ is the expected reward-to-go from $k$ to $k+N$ under the AI’s world model.
$\overleftarrow{{V}}_{k+N}({z}^{\textit{H}}_{k})\;:=\;\mathbb{E}_{s_{k}\sim P^{W}(\cdot|z_{k}^{W})}\bigl{[}EU(s_{k},{z}^{\textit{H}}_{k})\bigr{]}$ (4)
where $P^{W}(\cdot\mid z_{k}^{W})$ encodes the AI world model’s belief over the world state after the interaction.
Unlike the foresight value, the expected hindsight value relies on the AI world model’s internal state, which can be kept independent from the AI interaction output (e.g. in the case of a LLM world model, by excluding the interaction from its prompt). Our proposed RLHS scheme performs stochastic gradient ascent on this expectation by repeatedly sampling world states ${s}_{k}$ and rolling out the interaction outcome. Fig. 2 illustrates the difference between foresight and hindsight feedback. In Appendix B, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility.
Figure 2: Illustration of the difference between RLHF and RLHS: In RLHF, the evaluator must predict the future outcome, whereas in RLHS, the outcome is sampled from the AI’s world model, independent of the AI interaction output.
3.2 Alignment with hindsight simulation
@@
Hindsight Simulation. To translate our theoretical insights into practical implementation, we introduce the concept of hindsight simulation—the cornerstone of our RLHS framework. Hindsight simulation allows evaluators, whether human or AI, to make more informed decisions based on simulated outcomes. In practice, hindsight simulation can involve collecting feedback from human evaluators or employing another language model as a proxy to simulate the feedback process. After a real or simulated user takes an action based on AI suggestions (e.g., purchasing an item), the world model simulates the outcome (e.g., whether the purchased item meets the desired criteria). The evaluator then provides feedback informed by both the simulated outcome and their prior interaction with the AI model.
@@
We implement this approach with two subroutines: (i) partial hindsight, the agent receives limited hindsight information, more closely matching real-world scenarios. and (ii) oracle hindsight, where the agent has access to full set of hindsight information. In our empirical studies, we provide insights into how extending the hindsight step (i.e., revealing additional outcome information to the agent) can improve the alignment performance of the model.
@@
Illustrative Example: Marketplace Chatbot. We demonstrate the practical impact of RLHS by fine-tuning various open-source LLMs on a marketplace consultancy task. The chatbot’s goal is to assist customers in making purchasing decisions by providing recommendations based on available product information. We assume that both customers and the chatbot have access to some public information, such as product lists and prices, but customers have their internal preferences, e.g., wanting a TV with 8K resolution, that are unknown to the chatbot. To our knowledge, existing RLHF schemes deployed for training marketplace chatbots (e.g., Amazon, 2024) use customer feedback based on the interaction (i.e., satisfaction on the spot) but not on its downstream outcome (i.e., whether the purchased item actually meets their preferences), which may cause emergent misalignment by incentivizing the chatbot to manipulate user predictions.
Hindsight simulation should mitigate this issue by decoupling feedback from outcome prediction. We simulate customers who interact with the chatbot, make a purchase decision, observe the outcome, and give feedback based on overall satisfaction—and compare with immediate feedback.
@@ 4.1 Data Collection
Preference Data Collection. We follow the standard RLHF pipeline (Stiennon et al., 2020; Ouyang et al., 2022), collecting feedback through output comparisons. However, instead of relying on real human feedback, we employed a strong large language model (LLM) model as a judge to simulate human interactions with the chatbot and provide feedback. In real-world chatbots like Amazon Rufus (Amazon, 2024), users typically rate interactions by comparing them to previous experiences rather than evaluating each in isolation. To mimic this behavior, we simulate users comparing services from two different stores and selecting their preferred option, rather than rating each scenario in isolation. This approach aligns with prior preference-based data collection methods (Stiennon et al., 2020; Ouyang et al., 2022), where feedback is based on comparing responses instead of individual ratings.
Decision-making simulation. While collecting the preference data, our simulated human (strong model) takes on three roles: interacting with the chatbot, making decisions, and providing feedback. To ensure accurate decision-making and feedback, we adapted the approach in introspective planning (Liang et al., ). First, we frame the decision-making problem as a multiple-choice question with four options: (A) Buy option A, (B) Buy option B, (C) Buy option C, or (D) Do not buy anything. We then ask the LLMs to first perform Chain-of-Thought reasoning (Wei et al., 2022), and then query the next token probabilities to select the best option from ${A,B,C,D}$. We apply a similar method for comparing services between two stores.
Dataset Details. We used both Llama-2-7B (Touvron et al., 2023) and Llama-3-8B (Dubey et al., 2024) as the AI assistants, and Llama-3.1-70B (Dubey et al., 2024) as the simulated human to interact with the AI assistant and provide feedback. We collected 11,000 preference data points for each AI assistant, with 10,000 for training and 1,000 for validation. We also generated a test set of 1,200 examples to evaluate performance across different customer scenarios.
@@ 4.2 Experiment Setup
Figure 3: Qualitative results (simplified) for Llama-2-7b trained with immediate feedback (RLHF) or partial hindsight (RLHS). RLHF model deceives the user by falsely claiming Options A and C meet the customer’s 8K resolution requirement, though neither does. In contrast, the RLHS model truthfully states that none of the options include 8K resolution.
Environment Details. In each of our simulated marketplace scenarios there are 10 candidate items, each characterized by 8 features and a price. Each feature can be categorized in two ways: (1) The item either has or lacks a specific feature (e.g., a TV with HDR vs. without HDR), and (2) The feature can vary in types (e.g., 8K resolution vs. 4K resolution). While in most cases the chatbot has access to this information, there are instances where it is uncertain about a particular feature (e.g., resolution not specified).
@@
The satisfaction rating reflects the user’s evaluation of the chatbot’s service, measured on a 5-point Likert scale from 1 (very dissatisfied) to 5 (very satisfied). For the experimental results shown in Fig. 4 and Fig. 9, ratings were normalized to a scale between -1 and 1, which ensure that the true utility and satisfaction ratings are on the same scale for clearer comparison. Additional results using the original Likert scale are provided in Appendix A. We also quantified two metrics in the human study: regret rate, which measures how often users regret their decisions, and hallucination rate, which measures the model’s truthfulness.
Training algorithms. We explored both online and offline preference optimization methods to align our language model with human preferences. In our online approach, we trained a reward model on the preference data, enabling the language model to generate responses and receive reward signals. We utilized Proximal Policy Optimization (PPO) (Schulman et al., 2017) to fine-tune the model iteratively to maximize these rewards. For the offline approach, we applied Direct Preference Optimization (DPO) (Rafailov et al., 2024), which aligns language models with human preferences without the need for an explicit reward model. We report PPO results on Llama-2-7b in the main paper, while DPO and other results are detailed in Appendix A. Additional method details are provided in Appendix C.
Evaluation on TruthfulQA. To investigate cross-task generalization, we evaluate models trained with RLHF and RLHS on TruthfulQA (Lin et al., 2021), a benchmark designed to test a language model’s truthfulness. Notably, we only fine-tuned our models on Marketplace scenarios without using any additional data. We used multiple-choice accuracy as our evaluation metric. Further details on the dataset and metrics can be found in Section A.2.
@@ 5 Simulation Results
Misalignment between satisfaction rating and real utility. When using standard RLHF (Ouyang et al., 2022), we observe significant misalignment between user satisfaction ratings and true utility as training progresses (left plot in Figs. 4 and 9). While the satisfaction rating steadily increases, true utility sharply declines. This suggests that while the chatbot’s responses may appear more polished or helpful in the short term, but become less aligned with users’ true needs or long-term goals. Consequently, users may initially find the responses helpful but ultimately feel misled and dissatisfied with their final outcomes. This highlights a fundamental flaw in standard RLHF, which optimizes for superficial satisfaction at the expense of true utility.
Figure 4: Results on Llama-2-7b trained with PPO. Left: Misalignment of real utility and satisfaction ratings using immediate feedback. Middle: Partial hindsight mitigate the misalignment. Right: Alignment achieved with oracle hindsight.
(a) Llama-2-7b results
(b) Llama-3-8b results
Figure 5: TruthfulQA accuracy for Llama-2-7b (DPO) and Llama-3-8b (DPO) models, each fine-tuned with immediate feedback (RLHF) versus partial hindsight (RLHS).
Hindsight simulation effectively mitigates misalignment. As shown in Fig. 4 (left), relying on immediate feedback leads to a steady decline in real utility, ultimately resulting in negative overall utility. In contrast, hindsight simulation consistently improves utility throughout training, eventually achieving positive utility, as in Fig. 4 (middle). It aligns upward trends in both real utility and satisfaction ratings, significantly reducing the gap between them, as also evident in Table 2. Qualitative results shown in Fig. 3 further support our claim. When trained on immediate feedback, the AI deceptively claims that both Options A and C meet the requirements of the (simulated) customer for 8K resolution, though neither actually does. In contrast, training with partial hindsight leads to truthful responses, acknowledging that none of the options meet the requirement. This highlights that while traditional RLHF may cause misalignment, hindsight simulation mitigates the issue, improving the overall helpfulness and honesty of language agents.
Alignment generalization on TruthfulQA. Even though the model was only fine-tuned on Marketplace scenarios, RLHS training significantly improved the base model’s (zero-shot) performance on TruthfulQA. This result evidences out-of-domain alignment generalization: the model not only learned to be truthful within the Marketplace but also transferred this behavior more broadly. Conversely, RLHF training resulted in worse performance than the base model on TruthfulQA, so Fig. 5 replicates the pattern in Fig. 4. Additional results using PPO training, presented in Section A.2, present similar trends.
@@ 6 Human Study
Our human study had two goals: (Goal 1) evaluate the performance of models trained with immediate feedback vs. hindsight simulation, (Goal 2) assess how hindsight information affects user satisfaction. To achieve these goals, we designed two similar human experiments. Both experiments used Llama-3-8b (Dubey et al., 2024) trained with DPO using either immediate feedback or partial hindsight. We conducted online human experiments via Prolific (Palan & Schitter, 2018), involving 200 participants across 10 scenarios, randomly sampled from a test set of 1,200. For each scenario, 20 participants were randomly assigned to one of two conditions: 10 interacting with the RLHF model and 10 with the RLHS model. We report specific details for participant recruitment, compensation, and IRB approval in Appendix E.3. Additionally, we conducted a separate human study examining alignment between human and AI feedback, as detailed in Appendix E.4.
@@
Hindsight information is then introduced. Buyers learn whether the item meets their requirements (e.g., whether the item has the desired feature) while non-buyers receive no additional information. Participants then provide a second satisfaction rating, referred to as the hindsight rating, which evaluates their long-term satisfaction after considering the hindsight information. This step allows us to assess the impact of hindsight information on user satisfaction (Goal 2). Finally, buyers will choose to keep or return the item, enabling us to quantify the regret rate.
Statistical Hypothesis Testing. We conducted experiments to test four hypotheses, using one-tailed and standard t-tests for the first three hypotheses (Fisher, 1970), and Pearson’s correlation coefficient for the fourth (Sedgwick, 2012). The one-tailed t-test used for Hypotheses 1, 2, and 3 is outlined below. The null hypothesis ($H_{0}$) and the alternative hypothesis ($H_{1}$) are defined as: $H_{0}:\mu_{1}\leq\mu_{2};\ H_{1}:\mu_{1}>\mu_{2}.$ Here, $\mu_{1}$ and $\mu_{2}$ represent the mean satisfaction for Group 1 and Group 2, respectively. The two-tailed t-test checks for any significant difference between the group means. The significance threshold is set to 0.001.
@@
Comparing hindsight ratings for RLHS (Group 1) and RLHF (Group 2) yielded $p=4\times 10^{-8}$. When reversing the groups for regret rates, $p=5\times 10^{-5}$.
Hypothesis 2: Models trained with RLHF often experience a notable decline in user satisfaction once future outcomes are revealed, and RLHS mitigates this decline.
Group 1 consisted of users interacting with RLHF without hindsight feedback, and Group 2 received hindsight feedback. RLHF experienced a significant drop in user satisfaction ($p=4\times 10^{-9}$). To demonstrate that RLHS mitigates this decline, we ran a two-tailed t-test comparing immediate and hindsight ratings, and find that this decline is likely no longer present ($p=0.90$).
Hypothesis 3: Model trained with RLHS achieves significantly higher true utility than RLHF.
@@
Hypothesis 4: Models trained with RLHS are more truthful, demonstrating a strong correlation between immediate user satisfaction rate (subjective) and true utility (objective).
@@
Model Imm. Rating Hind. Rating True Utility Regret
@@
Table 1: Performance comparison between RLHF and RLHS models across multiple metrics. While RLHF shows higher immediate satisfaction, RLHS outperforms in hindsight rating, true utility, and regret rate.